# Fast rotations in galaxies at cosmic noon indicate central concentration of stars, dark matter or massive black holes

### v1

@@ 4.1 Surface Brightness best-fit models
In Figures 2, 3 and 4, we show the best-fit models for the gas and stellar surface brightness of ID1, ID3 and ID13, respectively. The corresponding corner plots are shown in Appendix A. We report in Table 4 the parameters defining the surface brightness for the gas disc, the stellar disc and the stellar bulge: the effective radius (), the centre ( and ), the ellipticity () and position angle (PA). Under the thin disc assumption, the ellipticity values of the stellar disc result in values of the inclination angles consistent with those used in rizzo23 to fit the gas kinematics. In Table 4, we also report the derived bulge-to-total luminosity ratio (B/T) for the stars.
@@ 4.2.1 Fiducial mass models
We show the results of the fiducial models for the three galaxies in Figure 5 and in Table 5. The cornerplots for these fits are shown in Appendix C. As discussed in Section 3.3.1, we tried to constrain the three main mass components: the total stellar mass, the total gas mass and the dark matter halo virial mass. Even though we have estimates of the stellar masses from SED fitting, the stellar masses might have systematic uncertainties of 0.3 dex (e.g. pacifici23; Choe_2025) and the possible differences in due to different stellar populations in the bulge and disc are unknown. Moreover, the mass contribution of the bulge relative to the disc is very uncertain, considering that for the three targets, the bulges seem to be unresolved (see Section 4.1). For the gas component, the emission line ratios and are uncertain for the CO emission of ID1 and ID3. For ID13, we have a good estimate of the [CI] line ratio by assuming a typical gas excitation temperature, however, the typical values of from the literature tend to be too high for the observed rotation curve, as discussed in Section 3.3.1.
@@
Similarly to the results in romanoliveira24, we found that the total stellar masses are generally well constrained with small uncertainties, as seen in the corner plots (see Figures 15 – 17). This is due to the fact that the stars dominate the gravitational potential in the inner region of the galaxies, and more specifically, the stellar bulge is the main contributor to this. The total stellar masses derived from our fiducial fits are significantly different from those reported in Table 1, which were obtained from SED fitting, if the quoted errors are fair representations of the uncertainties. However, if we consider systematic errors of the order of 0.3 dex for the SED models (see pacifici23; Choe_2025), then the masses obtained for the three galaxies are borderline within the uncertainties. For ID1 in particular, the dynamical stellar mass is significantly higher than that expected from the SED fitting; however, the SED fitting used was obtained before the JWST data became available. A preliminary SED fitting, including the JWST bands, skews the SED mass to 2-3 times higher and closer to our dynamical estimate.
@@
Even though we do not obtain precise constraints for the , the median values we found do not differ significantly from what was expected. If we divide the shown in Table 5 by the expected line ratios from the literature (0.83, 0.11 and 0.57 for ID1, ID3 and ID13 respectively), the estimated values are 3.7, 9.1 and 1.8 M. For ID1 and ID13, these values are within the range of expected values in the literature; albeit, ID13 is on the lower end. For ID3, the is quite high, but it might be more indicative that the line ratio of 0.11 for the CO (5-4)/CO (1-0) transitions is inappropriate, as the estimate for the line ratio of CO (5-4) is based on very few objects (see Section 3.3.1). If we divide the measured of ID3 by the expected M for starburst galaxies, we find that the luminosity line ratio would be 1.25. Alternatively, if the is closer to 4.0 M, then the line ratio would be 0.25. The latter might not be unreasonable for a starburst galaxy, as in dunne22 they show that an of 4 M is a near-universal average value for metal-rich galaxies. However, it is challenging to break the degeneracy between and using the constraints in hand.
To summarise, in Figure 6 we show an overall comparison of the stellar and gas masses obtained from our fiducial mass models and the expected values from the literature. For the stars, we compare our fiducial stellar masses with the stellar masses and their uncertainties obtained from SED models given in Table 1. For the gas, we compare our fiducial gas mass with an estimate of the gas mass using Equation 15 and the CO/[CI] luminosities reported in Table 1. These gas mass estimates are obtained using typical values in the literature for of 3.6, 0.8 and 7 M and line ratios of 0.83, 0.11, and 0.57 for ID1, ID3, and ID13, respectively. To estimate the errorbars shown, we assume a range of possible values for the : for ID1, as a main-sequence galaxy observed in a CO transition, we assume that the can vary between 3 and 4 M; for ID3, a starburst galaxy observed in a CO transition, we assume that the can vary between 0.8 and 4 M; for ID13 we assume that the ranges between 1 and 10 M (see discussion in Section 3.3.1). For the errors regarding the line ratios, we use the errors reported in boogaard20 for ID1 and ID3 and we assume no errors in the line ratio of ID13.
@@ 4.2.2 Models with the baryonic parameters fixed
For the gas masses, as mentioned in Section 3.3.1, we assume that the CO luminosity-to-H2 mass conversion factor () is 3.6 and 0.8 M for the main-sequence galaxy ID1 and for the starburst galaxy ID3, respectively (daddi10a; kirkpatrick19). We also assume that the CO luminosity line ratios from boogaard20 of 0.83 for the luminosity line ratio of CO (2-1)/CO (1-0) for ID1 and 0.11 for the line ratio of CO (5-4)/CO (1-0) for ID3.
On the other hand, the gas emission of ID13 is traced with the [CI] (3P2 - 3P1) emission, which presents some caveats that are explained in Section 3.3.1. The typical values found for the [CI] luminosity to gas mass conversion factor in the literature tend to be around 7 M; however, they can range from 1 to 10 M. If we assume a gas excitation temperature of as suggested by previous works (e.g. walter11; valentino20b; henriquezbrocal22), we obtain a line ratio (L’/L’) of 0.57. This gives us a total molecular gas of log Mgas/M⊙ = 10.6 when assuming the typical value of = 7 M, or a total molecular gas of log Mgas/M⊙ = 9.7 when assuming a conservative = 1 at the lower end of the expected values. Translating this to the velocity contribution to the rotation curve, it would mean that the gas disc would contribute at (the last point of the rotation curve) with or for of 7 and 1 M, respectively. The former value is physically impossible, as the observed rotation at is , and it needs to accommodate both the gas disc and the stellar disc, contributing at least when considering the stellar mass of ID13 reported in Table 1 and the bulge-to-total luminosity ratio in Table 4. Therefore, we kept the conservative M to fix the gas mass of ID13.
For the stellar masses, we use the values obtained from SED models for ID1 and ID3 reported in rizzo23 and shown in Table 1. For ID13, we assume a more conservative stellar mass: we take the SED fit stellar mass from Table 1 and subtract 0.3 dex, which is the suggested value of systematic uncertainties in SED models (pacifici23). We made this choice because, if we use the stellar mass in Table 1, the contributions from the stellar bulge and disc alone already overpredict the rotation in the galaxy. In all cases, we kept the assumption of a difference in of 1.4 between the bulge and disc components, and we maintained the shape of the surface density of the gas disc, stellar disc, and stellar bulge as obtained from the surface brightness fitting and already used in the fiducial models. We thus obtain mass models in which the baryonic components are completely fixed and the parameters for the NFW dark matter halo (i.e., and concentration) are allowed to vary.
@@
In these models, we use a flat log prior for the NFW halo mass (M200) that range, from 10.5 to 15 and a flat prior limited from 2.5 to 20 for the NFW halo concentration (). The concentration prior range roughly represents the possible concentrations expected for NFW dark matter halos in the local Universe and up to with virial masses between 1011 and 1015 M⊙ (dutton14; ludlow14).
We show these new models in Figure 7 and present the results in Table 6, where it is immediately apparent that, for ID1 and ID3, the models struggle to reproduce the observed rotation curve in a manner similar to that of the fiducial models. The main reason for the results for ID1 and ID3 is that the almost flat shape of the observed rotation curves cannot be reproduced by an NFW halo with a concentration value expected in CDM cosmology. We experimented by leaving the concentration free with a much larger prior range and we found that the best model tends towards concentration values upwards of 50, above a 5- level of statistical significance from the expected concentration-mass relation in dutton14. Unlike ID1 and ID3, which tend to be dark matter dominated in this fit, ID13 is dominated by the stellar component. We note that this fit is very similar to the fiducial one, given that we are assuming a stellar mass similar to that found with the fiducial models. This way, the only change occurs by keeping the gas mass fixed and allowing the dark matter halo mass to vary freely. This, however, leads to a small dark matter halo mass, as we do not impose the cosmological baryonic fraction, while the dark matter halo concentration has large uncertainties of . We note, again, that this fit only works when assuming conservative contributions of both the stellar and gas masses.
@@ 5.1.1 Sensitivity of the kinematic models to bulge contribution
When we assume that the bulge and stellar disc have the same mass-to-light ratio (; dashed purple line in Fig. 8), the resulting model underestimates the observed rotation velocity at the innermost data point. To investigate whether this discrepancy can be mitigated, we constructed kinematic models, i.e. mock emission line datacubes. Essentially we took the the kinematic parameters defining the best-fit models in rizzo23 and changed the first points of the rotation curves, assigning the values we obtained from the above dynamical models with . These values are , , and for IDs 1, 3, and 13, respectively. We then computed the values for each of these models and compared them to the best-fit values from rizzo23 using the Akaike Information Criterion (AIC). This comparison indicates that, for ID1, the two models are not significantly different, suggesting that the first data point may be biased toward higher velocities. In contrast, for IDs 3 and 13, the AIC differences are substantially larger ( 1000), implying that the fixed models provide a significantly poorer fit compared to the best-fit models reported in rizzo23.
@@ 5.1.2 Massive bulges
In this Section, we explore potential reasons for the need for a different mass-to-light ratio for the bulge and the disc. At , assuming different mass-to-light ratios for the bulge and the disc is standard (lelli16b). This assumption is based on the expected difference in the age/metallicity of the older stellar populations in the bulge, which can drive a difference in colours and, therefore, in the mass-to-light ratios. If the bulges in the ALPAKA galaxies are sufficiently old compared to the disc, then the difference could be similar to those of local galaxies where the is a common assumption based on stellar population synthesis models (schombert14; lelli16b).
The main limitation involved in the analysis of the difference in in the ALPAKA galaxies is the challenge of obtaining constraints on the star formation history. However, considering the age of the Universe at the three redshifts of these galaxies, we can estimate the maximum difference between the mass-to-light ratios of the bulge and disc. schombert19 shows that the difference in between a 3 Gyr old stellar population with solar metallicity and a younger population of 300 Myr with metallicity of [Fe/H] = can reach 1.6. Therefore, based on these values, we test a scenario in which the bulge and the disc have a difference of 1.6, to represent the extreme scenario of an old bulge and a constant star forming disc. This is not a fit, but only a rescaling of the bulge and disc contributions. We show this in the top row of Figure 8, where we compare the fiducial model with . We see that in all cases, the higher difference improves the fit in the inner region, as it tends to boost the central contribution of the stars. Whether values of above one are realistic for the stellar populations of these galaxies remains, however, an open question.
Another way of viewing the needed difference in the mass-to-light ratios of the bulge and the disc is that, instead of being driven by the stellar populations, it is driven by dust obscuration. Despite these galaxies being observed in the rest-frame near infrared, the central light of the galaxies may still be significantly attenuated (e.g., polletta24), which leads to two distinct effects: i) determining the structure of the bulge (namely the effective radius) may be difficult, and ii) the total intensity of the bulge may be underestimated. Both of these could explain why the stellar surface brightness fits tended to underestimate the bulge contribution (see Section 4.1). To quantify the effect of dust attenuation, we assumed a total-to-selective extinction ratio of 4.05 and a reddening curve (calzetti00). The relation between observed and intrinsic luminosity is given by
@@
Following this, we found that the three ALPAKA bulges must have, on average, extinctions of AV of 1.7, 3.0 and 1.7 mag above that of the disc, respectively, for ID1, ID3, and ID13. We note that these results can vary under different assumptions of the dust model, as they can depend on the spectral types and physical properties of the galaxies (kriek13). However, if we assume that the dust is mostly concentrated close to the bulge, then an inner AV of 2 mag is in line with typical dusty galaxies at z (e.g. dacunha15). Moreover, some dusty star forming galaxies at observed with JWST have been found to have strong colour gradients and heavily obscured bulges (miller22; hodge24; polletta24). We conclude that considering a larger value for the mass-to-light ratio of the bulge with respect to that of the disc can be a realistic choice.
@@ 5.1.3 Alternative to massive bulges as main contributors to the inner rotation velocities
Another way to boost the rotation velocity in the innermost region is to consider the presence of a supermassive black hole. Despite the fact that we use data that have not been taken to resolve the sphere of influence, the addition of a SMBH can potentially contribute to the inner rotation curve (e.g., ruffa23; zhang24). In the second row of Figure 8, we show the effect of a SMBH on the mass models by comparing a black hole that follows the black hole–stellar-mass relation for local galaxies from reines15 and the relation for galaxies at from pacucci23. To explore this scenario, we compare the mass models including SMBHs with models similar to our fiducial models, but without the assumption of the difference between bulge and disc; we call this the baseline model. We note that we are not refitting the rotation curves; rather, we are only adding a black hole component to the best-fit rotation curve.
@@
We found that, in all three cases, a supermassive black hole that follows the local relation from reines15 does not produce results that are significantly different from the baseline model. On the other hand, when using the relation from pacucci23, which implies a black hole mass of approximately 1.5% of the total stellar mass instead of only 0.25% as in reines15, we find a significant increase in the central velocity of the model and an overall change in the shape of the rotation curve. This is enough to bring the central velocity within the observational uncertainties for ID1; however, it is still not sufficient to reproduce the inner rotation in the other two galaxies. If we were to bring the mass models to overlap with the central observed velocity at a level, then the black hole velocity contributions to the inner points should be , 150, , which translate to SMBH masses of 1.5%, 3.0%, and 4.4% of the total stellar mass for ID1, ID3 and ID13, respectively.
@@
The high black hole to stellar mass ratios of high- galaxies are an ongoing debate in the literature, as many new JWST observations suggest the presence of overly massive black holes when compared to the local relation (e.g., goulding23; furtak24). Although the presence of a black hole does not seem to be enough to reproduce the rotation curves of our targets, our results are promising from another perspective, as they show that with a resolution 3 to 4 times better than the current data, we could attempt to constrain SMBH masses at .
@@
Lastly, another possibility is the existence of a radial gradient in the CO luminosity to gas mass conversion factor, which has been suggested in previous studies of nearby galaxies (e.g., sandstrom13; bigiel20). This is not straightforward to implement on the mass decomposition models, as the assumption that the gas is distributed in an exponential disc is no longer valid if we apply a radial variation of the .
@@
However, we can conduct a simple experiment in which we derive the mass density of the galaxies with a radially varying and estimate the circular velocity at each radius from the enclosed mass, V. We note that this implicitly assumes that the mass is distributed in a spherical geometry; however, the circular velocity of a mass component distributed in a sphere is at most 20% different from its circular velocity as if it were distributed in a disc (binney08).
@@
In order for the gas mass distribution to be higher in the centre than in the outskirts, the factor gradient needs to be negative. The factor depends on the physical conditions of the gas (e.g., density, temperature and gas-phase metallicity) and it usually has a positive radial gradient in the galaxies where it has been observed (e.g., sandstrom13; bigiel20). This is because the is lower in high-metallicity regions, which are usually found in the centre of galaxies. For this gradient to explain the inner velocity discrepancy, the galaxies should have a large positive metallicity gradient. This is uncommon in the local Universe; however, at high-, such gradients are possible. For instance, some works found positive and flat metallicity gradients at z 1 (jones13; troncoso14; leethochawalit16; wang22; venturi24). To maximise the gas contribution in the central regions, we assume the metallicity gradient of the galaxy YD4, part of the proto-cluster 2744-z7p9OD at z 7.9, which was reported in venturi24 to have a metallicity gradient () of dex kpc-1. This is the best example of a positive metallicity gradient in a high- galaxy, albeit troncoso14 also reports similar gradients in galaxies at z=. We then consider an upper-limit metallicity gradient of +0.3 dex kpc-1, corresponding to 1- above the median value measured for YD4, and compute for each galaxy the associated change in . Using the relation between and metallicity from sandstrom13, we convert this metallicity gradient into a radial gradient. This yields , which we adopt as the slope of the radial profile. For the intercept, we assumed that the of the inner point of the rotation curve to be the same as the baseline, so we can compare the change of the gas circular velocity with the baseline model.

### v2

@@ 4.1 Surface Brightness best-fit models
In Figures 2, 3 and 4, we show the best-fit models for the gas and stellar surface brightness of ID1, ID3 and ID13, respectively. The corresponding corner plots are shown in Appendix A. We report in Table 4 the parameters defining the surface brightness for the gas disc, the stellar disc and the stellar bulge: the effective radius (), the centre ( and ), the ellipticity () and position angle (PA). Under the thin disc assumption, the ellipticity values of the stellar disc result in values of the inclination angles consistent with those used in Rizzo et al. (2023) to fit the gas kinematics. In Table 4, we also report the derived bulge-to-total luminosity ratio (B/T) for the stars.
@@ 4.2.1 Fiducial mass models
We show the results of the fiducial models for the three galaxies in Figure 5 and in Table 5. The cornerplots for these fits are shown in Appendix C. As discussed in Section 3.3.1, we tried to constrain the three main mass components: the total stellar mass, the total gas mass and the dark matter halo virial mass. Even though we have estimates of the stellar masses from SED fitting, the stellar masses might have systematic uncertainties of 0.3 dex (e.g. Pacifici et al., 2023; Choe et al., 2025) and the possible differences in due to different stellar populations in the bulge and disc are unknown. Moreover, the mass contribution of the bulge relative to the disc is very uncertain, considering that for the three targets, the bulges seem to be unresolved (see Section 4.1). For the gas component, the emission line ratios and are uncertain for the CO emission of ID1 and ID3. For ID13, we have a good estimate of the [CI] line ratio by assuming a typical gas excitation temperature, however, the typical values of from the literature tend to be too high for the observed rotation curve, as discussed in Section 3.3.1.
@@
Similarly to the results in Roman-Oliveira et al. (2024), we found that the total stellar masses are generally well constrained with small uncertainties, as seen in the corner plots (see Figures 15 – 17). This is due to the fact that the stars dominate the gravitational potential in the inner region of the galaxies, and more specifically, the stellar bulge is the main contributor to this. The total stellar masses derived from our fiducial fits are significantly different from those reported in Table 1, which were obtained from SED fitting, if the quoted errors are fair representations of the uncertainties. However, if we consider systematic errors of the order of 0.3 dex for the SED models (see Pacifici et al., 2023; Choe et al., 2025), then the masses obtained for the three galaxies are borderline within the uncertainties. For ID1 in particular, the dynamical stellar mass is significantly higher than that expected from the SED fitting; however, the SED fitting used was obtained before the JWST data became available. A preliminary SED fitting, including the JWST bands, skews the SED mass to 2-3 times higher and closer to our dynamical estimate.
@@
Even though we do not obtain precise constraints for the , the median values we found do not differ significantly from what was expected. If we divide the shown in Table 5 by the expected line ratios from the literature (0.83, 0.11 and 0.57 for ID1, ID3 and ID13 respectively), the estimated values are 3.7, 9.1 and 1.8 M. For ID1 and ID13, these values are within the range of expected values in the literature; albeit, ID13 is on the lower end. For ID3, the is quite high, but it might be more indicative that the line ratio of 0.11 for the CO (5-4)/CO (1-0) transitions is inappropriate, as the estimate for the line ratio of CO (5-4) is based on very few objects (see Section 3.3.1). If we divide the measured of ID3 by the expected M for starburst galaxies, we find that the luminosity line ratio would be 1.25. Alternatively, if the is closer to 4.0 M, then the line ratio would be 0.25. The latter might not be unreasonable for a starburst galaxy, as in Dunne et al. (2022) they show that an of 4 M is a near-universal average value for metal-rich galaxies. However, it is challenging to break the degeneracy between and using the constraints in hand.
To summarise, in Figure 6 we show an overall comparison of the stellar and gas masses obtained from our fiducial mass models and the expected values from the literature. For the stars, we compare our fiducial stellar masses with the stellar masses and their uncertainties obtained from SED models given in Table 1. For the gas, we compare our fiducial gas mass with an estimate of the gas mass using Equation 15 and the CO/[CI] luminosities reported in Table 1. These gas mass estimates are obtained using typical values in the literature for of 3.6, 0.8 and 7 M and line ratios of 0.83, 0.11, and 0.57 for ID1, ID3, and ID13, respectively. To estimate the errorbars shown, we assume a range of possible values for the : for ID1, as a main-sequence galaxy observed in a CO transition, we assume that the can vary between 3 and 4 M; for ID3, a starburst galaxy observed in a CO transition, we assume that the can vary between 0.8 and 4 M; for ID13 we assume that the ranges between 1 and 10 M (see discussion in Section 3.3.1). For the errors regarding the line ratios, we use the errors reported in Boogaard et al. (2020) for ID1 and ID3 and we assume no errors in the line ratio of ID13.
@@ 4.2.2 Models with the baryonic parameters fixed
For the gas masses, as mentioned in Section 3.3.1, we assume that the CO luminosity-to-H2 mass conversion factor () is 3.6 and 0.8 M for the main-sequence galaxy ID1 and for the starburst galaxy ID3, respectively (Daddi et al., 2010; Kirkpatrick et al., 2019). We also assume that the CO luminosity line ratios from Boogaard et al. (2020) of 0.83 for the luminosity line ratio of CO (2-1)/CO (1-0) for ID1 and 0.11 for the line ratio of CO (5-4)/CO (1-0) for ID3.
On the other hand, the gas emission of ID13 is traced with the [CI] (3P2 - 3P1) emission, which presents some caveats that are explained in Section 3.3.1. The typical values found for the [CI] luminosity to gas mass conversion factor in the literature tend to be around 7 M; however, they can range from 1 to 10 M. If we assume a gas excitation temperature of as suggested by previous works (e.g. Walter et al., 2011; Valentino et al., 2020; Henríquez-Brocal et al., 2022), we obtain a line ratio (L’/L’) of 0.57. This gives us a total molecular gas of log Mgas/M⊙ = 10.6 when assuming the typical value of = 7 M, or a total molecular gas of log Mgas/M⊙ = 9.7 when assuming a conservative = 1 at the lower end of the expected values. Translating this to the velocity contribution to the rotation curve, it would mean that the gas disc would contribute at (the last point of the rotation curve) with or for of 7 and 1 M, respectively. The former value is physically impossible, as the observed rotation at is , and it needs to accommodate both the gas disc and the stellar disc, contributing at least when considering the stellar mass of ID13 reported in Table 1 and the bulge-to-total luminosity ratio in Table 4. Therefore, we kept the conservative M to fix the gas mass of ID13.
For the stellar masses, we use the values obtained from SED models for ID1 and ID3 reported in Rizzo et al. (2023) and shown in Table 1. For ID13, we assume a more conservative stellar mass: we take the SED fit stellar mass from Table 1 and subtract 0.3 dex, which is the suggested value of systematic uncertainties in SED models (Pacifici et al., 2023). We made this choice because, if we use the stellar mass in Table 1, the contributions from the stellar bulge and disc alone already overpredict the rotation in the galaxy. In all cases, we kept the assumption of a difference in of 1.4 between the bulge and disc components, and we maintained the shape of the surface density of the gas disc, stellar disc, and stellar bulge as obtained from the surface brightness fitting and already used in the fiducial models. We thus obtain mass models in which the baryonic components are completely fixed and the parameters for the NFW dark matter halo (i.e., and concentration) are allowed to vary.
@@
In these models, we use a flat log prior for the NFW halo mass (M200) that range, from 10.5 to 15 and a flat prior limited from 2.5 to 20 for the NFW halo concentration (). The concentration prior range roughly represents the possible concentrations expected for NFW dark matter halos in the local Universe and up to with virial masses between 1011 and 1015 M⊙ (Dutton and Macciò, 2014; Ludlow et al., 2014).
We show these new models in Figure 7 and present the results in Table 6, where it is immediately apparent that, for ID1 and ID3, the models struggle to reproduce the observed rotation curve in a manner similar to that of the fiducial models. The main reason for the results for ID1 and ID3 is that the almost flat shape of the observed rotation curves cannot be reproduced by an NFW halo with a concentration value expected in CDM cosmology. We experimented by leaving the concentration free with a much larger prior range and we found that the best model tends towards concentration values upwards of 50, above a 5- level of statistical significance from the expected concentration-mass relation in Dutton and Macciò (2014). Unlike ID1 and ID3, which tend to be dark matter dominated in this fit, ID13 is dominated by the stellar component. We note that this fit is very similar to the fiducial one, given that we are assuming a stellar mass similar to that found with the fiducial models. This way, the only change occurs by keeping the gas mass fixed and allowing the dark matter halo mass to vary freely. This, however, leads to a small dark matter halo mass, as we do not impose the cosmological baryonic fraction, while the dark matter halo concentration has large uncertainties of . We note, again, that this fit only works when assuming conservative contributions of both the stellar and gas masses.
@@ 5.1.1 Sensitivity of the kinematic models to bulge contribution
When we assume that the bulge and stellar disc have the same mass-to-light ratio (; dashed purple line in Fig. 8), the resulting model underestimates the observed rotation velocity at the innermost data point. To investigate whether this discrepancy can be mitigated, we constructed kinematic models, i.e. mock emission line datacubes. Essentially we took the the kinematic parameters defining the best-fit models in Rizzo et al. (2023) and changed the first points of the rotation curves, assigning the values we obtained from the above dynamical models with . These values are , , and for IDs 1, 3, and 13, respectively. We then computed the values for each of these models and compared them to the best-fit values from Rizzo et al. (2023) using the Akaike Information Criterion (AIC). This comparison indicates that, for ID1, the two models are not significantly different, suggesting that the first data point may be biased toward higher velocities. In contrast, for IDs 3 and 13, the AIC differences are substantially larger ( 1000), implying that the fixed models provide a significantly poorer fit compared to the best-fit models reported in Rizzo et al. (2023).
@@ 5.1.2 Massive bulges
In this Section, we explore potential reasons for the need for a different mass-to-light ratio for the bulge and the disc. At , assuming different mass-to-light ratios for the bulge and the disc is standard (Lelli et al., 2016a). This assumption is based on the expected difference in the age/metallicity of the older stellar populations in the bulge, which can drive a difference in colours and, therefore, in the mass-to-light ratios. If the bulges in the ALPAKA galaxies are sufficiently old compared to the disc, then the difference could be similar to those of local galaxies where the is a common assumption based on stellar population synthesis models (Schombert and McGaugh, 2014; Lelli et al., 2016a).
The main limitation involved in the analysis of the difference in in the ALPAKA galaxies is the challenge of obtaining constraints on the star formation history. However, considering the age of the Universe at the three redshifts of these galaxies, we can estimate the maximum difference between the mass-to-light ratios of the bulge and disc. Schombert et al. (2019) shows that the difference in between a 3 Gyr old stellar population with solar metallicity and a younger population of 300 Myr with metallicity of [Fe/H] = can reach 1.6. Therefore, based on these values, we test a scenario in which the bulge and the disc have a difference of 1.6, to represent the extreme scenario of an old bulge and a constant star forming disc. This is not a fit, but only a rescaling of the bulge and disc contributions. We show this in the top row of Figure 8, where we compare the fiducial model with . We see that in all cases, the higher difference improves the fit in the inner region, as it tends to boost the central contribution of the stars. Whether values of above one are realistic for the stellar populations of these galaxies remains, however, an open question.
Another way of viewing the needed difference in the mass-to-light ratios of the bulge and the disc is that, instead of being driven by the stellar populations, it is driven by dust obscuration. Despite these galaxies being observed in the rest-frame near infrared, the central light of the galaxies may still be significantly attenuated (e.g., Polletta et al., 2024), which leads to two distinct effects: i) determining the structure of the bulge (namely the effective radius) may be difficult, and ii) the total intensity of the bulge may be underestimated. Both of these could explain why the stellar surface brightness fits tended to underestimate the bulge contribution (see Section 4.1). To quantify the effect of dust attenuation, we assumed a total-to-selective extinction ratio of 4.05 and a reddening curve (Calzetti et al., 2000). The relation between observed and intrinsic luminosity is given by
@@
Following this, we found that the three ALPAKA bulges must have, on average, extinctions of AV of 1.7, 3.0 and 1.7 mag above that of the disc, respectively, for ID1, ID3, and ID13. We note that these results can vary under different assumptions of the dust model, as they can depend on the spectral types and physical properties of the galaxies (Kriek and Conroy, 2013). However, if we assume that the dust is mostly concentrated close to the bulge, then an inner AV of 2 mag is in line with typical dusty galaxies at z (e.g. da Cunha et al., 2015). Moreover, some dusty star forming galaxies at observed with JWST have been found to have strong colour gradients and heavily obscured bulges (Miller et al., 2022; Hodge et al., 2024; Polletta et al., 2024). We conclude that considering a larger value for the mass-to-light ratio of the bulge with respect to that of the disc can be a realistic choice.
@@ 5.1.3 Alternative to massive bulges as main contributors to the inner rotation velocities
Another way to boost the rotation velocity in the innermost region is to consider the presence of a supermassive black hole. Despite the fact that we use data that have not been taken to resolve the sphere of influence, the addition of a SMBH can potentially contribute to the inner rotation curve (e.g., Ruffa et al., 2023; Zhang et al., 2024). In the second row of Figure 8, we show the effect of a SMBH on the mass models by comparing a black hole that follows the black hole–stellar-mass relation for local galaxies from Reines and Volonteri (2015) and the relation for galaxies at from Pacucci et al. (2023). To explore this scenario, we compare the mass models including SMBHs with models similar to our fiducial models, but without the assumption of the difference between bulge and disc; we call this the baseline model. We note that we are not refitting the rotation curves; rather, we are only adding a black hole component to the best-fit rotation curve.
@@
We found that, in all three cases, a supermassive black hole that follows the local relation from Reines and Volonteri (2015) does not produce results that are significantly different from the baseline model. On the other hand, when using the relation from Pacucci et al. (2023), which implies a black hole mass of approximately 1.5% of the total stellar mass instead of only 0.25% as in Reines and Volonteri (2015), we find a significant increase in the central velocity of the model and an overall change in the shape of the rotation curve. This is enough to bring the central velocity within the observational uncertainties for ID1; however, it is still not sufficient to reproduce the inner rotation in the other two galaxies. If we were to bring the mass models to overlap with the central observed velocity at a level, then the black hole velocity contributions to the inner points should be , 150, , which translate to SMBH masses of 1.5%, 3.0%, and 4.4% of the total stellar mass for ID1, ID3 and ID13, respectively.
@@
The high black hole to stellar mass ratios of high- galaxies are an ongoing debate in the literature, as many new JWST observations suggest the presence of overly massive black holes when compared to the local relation (e.g., Goulding et al., 2023; Furtak et al., 2024). Although the presence of a black hole does not seem to be enough to reproduce the rotation curves of our targets, our results are promising from another perspective, as they show that with a resolution 3 to 4 times better than the current data, we could attempt to constrain SMBH masses at .
@@
Lastly, another possibility is the existence of a radial gradient in the CO luminosity to gas mass conversion factor, which has been suggested in previous studies of nearby galaxies (e.g., Sandstrom et al., 2013; Bigiel et al., 2020). This is not straightforward to implement on the mass decomposition models, as the assumption that the gas is distributed in an exponential disc is no longer valid if we apply a radial variation of the .
@@
However, we can conduct a simple experiment in which we derive the mass density of the galaxies with a radially varying and estimate the circular velocity at each radius from the enclosed mass, V. We note that this implicitly assumes that the mass is distributed in a spherical geometry; however, the circular velocity of a mass component distributed in a sphere is at most 20% different from its circular velocity as if it were distributed in a disc (Binney and Tremaine, 2008).
@@
In order for the gas mass distribution to be higher in the centre than in the outskirts, the factor gradient needs to be negative. The factor depends on the physical conditions of the gas (e.g., density, temperature and gas-phase metallicity) and it usually has a positive radial gradient in the galaxies where it has been observed (e.g., Sandstrom et al., 2013; Bigiel et al., 2020). This is because the is lower in high-metallicity regions, which are usually found in the centre of galaxies. For this gradient to explain the inner velocity discrepancy, the galaxies should have a large positive metallicity gradient. This is uncommon in the local Universe; however, at high-, such gradients are possible. For instance, some works found positive and flat metallicity gradients at z 1 (Jones et al., 2013; Troncoso et al., 2014; Leethochawalit et al., 2016; Wang et al., 2022; Venturi et al., 2024). To maximise the gas contribution in the central regions, we assume the metallicity gradient of the galaxy YD4, part of the proto-cluster 2744-z7p9OD at z 7.9, which was reported in Venturi et al. (2024) to have a metallicity gradient () of dex kpc-1. This is the best example of a positive metallicity gradient in a high- galaxy, albeit Troncoso et al. (2014) also reports similar gradients in galaxies at z=. We then consider an upper-limit metallicity gradient of +0.3 dex kpc-1, corresponding to 1- above the median value measured for YD4, and compute for each galaxy the associated change in . Using the relation between and metallicity from Sandstrom et al. (2013), we convert this metallicity gradient into a radial gradient. This yields , which we adopt as the slope of the radial profile. For the intercept, we assumed that the of the inner point of the rotation curve to be the same as the baseline, so we can compare the change of the gas circular velocity with the baseline model.